Supporting State-Sensitive Computation in a Dataflow System

Computation Structures Group Memo 294 March 1989

Paul S. Barth Rishiyur S. Nikhil

This report was supported in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-84-K-0099. Paul Barth was supported by a fellowship from Schlumberger Technology Corporation.